Findthe value of c that makes the expression a perfect square trinomial. Then write the expression as a square of a binomial. x^2 +3x + c
Answer by stanbon(75887)

Findthe value of c that makes the expression a perfect square trinomial. Then write the expression as a square of a binomial. x^2 +3x + c
-----
c = (3/2)^2 = 9/4
=====================
x^2+3x+(9/4)
= (x+(3/2))^2
=====================
Cheers,
Stan H.